a test of a driver's perception reaction time is being conducted on a special testing track with wet pavement and a driving speed of 50 kph. when the driver is sober, a stop can be made just in time to avoid hitting an object that is visible 40 m. ahead. after a few drinks of san miguel beer, under exactly the same condition, the driver fails to stop in time and strikes the object at a speed of 30 kph. determine the driver's perception-reaction time after he was drinking. assume coefficient of friction is 0.60.


1
Expert's answer
2021-01-19T02:31:23-0500

The driver's perception-reaction time is:


"t=t_1-t_2" ,


where "t_1" - the time from the moment, when the driver presses the brake, to the moment, when the car stops;

"t_2" - the time from the moment, when the driver presses the brake, to the moment, when the car strikes the object at a speed of 30 kmph.

First, find "t_1". The car speed depends on time as follows:


"v=v_0+at" ,


where "v_0" - the initial speed;

"a" - the acceleration.

In the first case "v=0", so


"t_1=-\frac{v_0}{a}" .


The acceleration can be found from the Newton's second law:


"ma=-\mu mg,"


"a=-\mu g."


Henсe, assuming that 50 kmph is 13,89 mps,


"t_1=\frac{v_0}{\mu g}=\frac{13.89}{0.6\cdot 9.81}=2.36\space s."


Then, find "t_2". In the second case "v=30\space kmph" or "8.33\space mps" so


"t_2=\frac{v_0-v}{\mu g}=\frac{13.89-8.33}{0.6\cdot 9.81}=0.95\space s."


"t=t_1-t_2=2.36-0.95=1.41\space s."


Answer: The driver's perception-reaction time is 1.41 seconds.